find the distance between (-3,-2) and (1,4)
a 27
b 2 square root 5
c 2 square root 13
d square root 10

Answer by edjones(4080)

-3-1=-4
-2-4=-6
(-4)^2+(-6)^2=16+36=52
sqrt(52)=2sqrt(13)
Ed